OpenStudy (anonymous):

You have been hired as contractors for your uncle's company. Your first day on the job you get a call from the Noether family who would like to install a pool in their backyard. Your job is to design the pool and provide some cost estimates.

OpenStudy (anonymous):

The pool should be a rectangular prism. Mrs. Noether wants the pool to hold between 1400 ft3 and 3200 ft3 of water and be from 6-10 feet deep. Other than those specifications, you are free to design the pool how you'd like! The estimated cost of the pool, besides labor, will come from the pool lining material that coats the inside of the pool and the amount of water needed to fill the pool. Determine the following: 1.The dimensions for the Noether pool that fit the above specifications 2. The amount of pool liner material that will be needed to completely line the inside of the pool (sides and bottom) 3. The amount of water that will be needed to fill the pool if the water needs to be 6 inches below the top of the pool *Please include the work involved in determining the above measurements or an explanation for how the measurements were determined. Draft a proposal for the Noethers that includes the above information along with a drawing of their pool. You may need to create more than one drawing to correctly convey your idea.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

What do you need help with?

OpenStudy (anonymous):

I do not understand how to do this problem. Like how to set it up, what to do

OpenStudy (whpalmer4):

Do you know what a rectangular prism looks like?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Okay. Draw one on a piece of paper and label the length, width, and height with l, w, h. We're going to have to figure out the values. We know that 6 <= h <= 10, and that 1400 <= l*w*h <= 3200 (given in the problem statement).

OpenStudy (whpalmer4):

Now, pick the depth you want the pool to be. It's a rectangular prism, so it will be a uniform depth from end to end.

OpenStudy (anonymous):

Ok 8. I basically need you to completely walk me through this question. If u can. I will medal

OpenStudy (whpalmer4):

Oh, boy, I'll do anything for a medal :rolleyes: Great, so we have a pool that is 8' deep. Now we need to find length and width so that 8*l*w is between 1400 and 3200 cu ft. What would you like the volume to be?

OpenStudy (anonymous):

1600 is good

OpenStudy (whpalmer4):

Okay, if the volume = 1600, and the height = 8, what is l*w?

OpenStudy (whpalmer4):

Remember, V = l*w*h

OpenStudy (anonymous):

1600/8=200/2

OpenStudy (whpalmer4):

1600/8 = 200, no?

OpenStudy (anonymous):

so 100 l and 100 h

OpenStudy (whpalmer4):

Uh, try V = 8*100*100 = ?

OpenStudy (anonymous):

that's 80,000

OpenStudy (whpalmer4):

is that between 1400 and 3200? :-)

OpenStudy (anonymous):

Ok so those are the dimensions, now about the pool liner

OpenStudy (whpalmer4):

no! answer my question...

OpenStudy (anonymous):

and no 80,000 is not in between those two integers

OpenStudy (whpalmer4):

okay, well, that's the volume of the pool you just proposed...

OpenStudy (anonymous):

oh so what numbers can we use besides 8. 8x?x?

OpenStudy (whpalmer4):

V = l*w*h We already chose h = 8, and V = 1600 1600 = l*w*8 Divide both sides by 8 200 = l*w Pick l and w so that that equation is satisfied...

OpenStudy (whpalmer4):

You did 200 = l + w

OpenStudy (anonymous):

8x100x4

OpenStudy (whpalmer4):

200 = 100*4?

OpenStudy (anonymous):

no these are new dimensions, could they work

OpenStudy (anonymous):

they are 3200

OpenStudy (whpalmer4):

You already chose V = 1600 and h = 8. That means 200 = l*w. If you choose 100 for l and 4 for w, does l*w = 200?

OpenStudy (anonymous):

40 and 5

OpenStudy (whpalmer4):

Okay. So we have a pool that is 8 feet deep, 40 feet long, 5 feet wide, and has a volume of 1600 cu ft. Now we need to work out the dimensions of the bottom and 4 sides so we can find their area (the amount of pool liner needed). Put in the dimensions on your diagram. You should have 3 differently sized rectangles — the bottom, the right and left side, and the front and back ends. Can you find the area of each rectangle? You'll add up the area of the bottom, the area of two side rectangles and the area of two end rectangles to get the total area of pool liner required.

OpenStudy (anonymous):

so 5 x 8 would be the first one?

OpenStudy (whpalmer4):

 What are the dimensions of the bottom?